Advanced Hoverboards can purchase a hoverboard manufacturing machine for $1 million that will last for 2 years. It will be able to produce and sell an average of 10,000 units per year at $97 net cash flow per unit. It expects that there is a 50% chance it will discover that the hoverboard is a successful product in a year. If it is a success, Advanced will expand production to 20,000 units for sale in year 2.

If it is not a success, Advanced will close up shop and sell its machine for after tax salvage value of $519841 thousand at the end of Year 1.

The discount rate is 13%.

What is the NPV of this project?

a. $821615

b. $848078

c. $1199920

d. $618059

e. $441789
Answer #1

NPV if success: 50% chance

Initial cost = -1000000

Cash flow in year 1 = 10000*97 = 970,000

Cash flow in year 2 = 20000*97 = 1940000

NPV= C0+ CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n

= -1000000+ 970000/1.13^1+ 1940000/1.13^2

=1377712

NPV if unsuccessful: 50% chance

Initial cost = -1000000

Cash flow in year 1 = 10000*97 + 519841= 1489841

NPV= C0+ CF1/(1+r)^1

= -1000000+ 1489841/1.13^1

=318443.4

Expected NPV = 1377712*50% + 318443.4*50%

= 848077.5

Correct option is b: 848078